What is the total pressure of a gas if the partial pressure of krypton is 730 mmHg, the partial pressure of neon is 1.50 atm and the partial pressure of helium is 45.6 mmHg?
1
Expert's answer
2020-04-03T11:51:48-0400

By Dalton's Law of Partial pressure,

The total pressure of a gas mixture is the sum of partial pressure of individual components of the gas mixture.

So, here total pressure of gas will be the sum of partial pressure of krypton,neon and helium.

"P=p_k+p_n+p_h=730 \ mm+" "1.5\ atm+45.6\ mm"

"=(730+1.5\times 760+45.6)\ mm=1915.6\ mm"